OpenStudy (rootbeer003):

What number should be added to the expression x2 + 14x to change it into a perfect square trinomial? 7 14 28 49

TheSmartOne (thesmartone):

what is half of 14?

OpenStudy (rootbeer003):

im guessing 49

OpenStudy (rootbeer003):

ok next

OpenStudy (anonymous):

bingo

TheSmartOne (thesmartone):

correcto :)

OpenStudy (rootbeer003):

Solve the following equation by transforming it into a perfect square trinomial. x^2 – 14x = 15 {–15, 29} {–1, 15} {57, 71} {–22, 36}

OpenStudy (anonymous):

don't be daft

OpenStudy (rootbeer003):

b?

OpenStudy (anonymous):

or we can do it that way if you like start with \[x^2-14+49=15+49\] go to \[(x-7)^2=64\]

OpenStudy (anonymous):

now there are two numbers whose square is 64, \(8\) and \(-8\) so \[x-7=8\\x=15\] or \[x-7=-8\\ x=-1\]

TheSmartOne (thesmartone):

bingo

OpenStudy (rootbeer003):

hold on im hungry

OpenStudy (anonymous):

that be the method of "completing the square" although you could have written \[x^2-14x-15=0\] then factored as \[(x-15)(x+1)=0\] etc

OpenStudy (rootbeer003):

okie im always right

OpenStudy (anonymous):

while you are up, get me a slider

OpenStudy (rootbeer003):

im not at taco bell

OpenStudy (anonymous):

so? i'm sure there is one close enough

OpenStudy (rootbeer003):

To solve the equation below by completing the square, what is your first step? 3x^2 + 18x = 21 Add 9 to each side of the equation. Add 81 to each side of the equation. Multiply both sides of the equation by 1/3 Multiply both sides of the equation by 1/9

OpenStudy (rootbeer003):

20 mins away

OpenStudy (anonymous):

i'll wait

OpenStudy (anonymous):

you have to have a leading coefficient of 1 to complete the square , so you have to divide all by 3

OpenStudy (anonymous):

which, for some inexplicable reason, they have written as "multiply by 1/3"

OpenStudy (anonymous):

time for a snack

OpenStudy (rootbeer003):

TIME FOR A SHNACK

TheSmartOne (thesmartone):

xD

OpenStudy (rootbeer003):

ok i back

TheSmartOne (thesmartone):

next question ;p

OpenStudy (rootbeer003):

oh its d

OpenStudy (rootbeer003):

i was in a trance

TheSmartOne (thesmartone):

it's not D

OpenStudy (anonymous):

one third, not one ninth

OpenStudy (rootbeer003):

oh sorry i meant one third

OpenStudy (rootbeer003):

its D for me

TheSmartOne (thesmartone):

it's C in the list you posted for us :P

OpenStudy (rootbeer003):

yup

OpenStudy (rootbeer003):

Use the quadratic formula to solve x^2 + 8x + 9 = 0. Estimate irrational solutions to the nearest tenth. {–3.7, 14.3} {–6.6, –1.4} {–18, 10} {–13.3, –2.7}

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x^2+%2B+8x+%2B+9+%3D+0 click on "approximate form"

OpenStudy (rootbeer003):

b

TheSmartOne (thesmartone):

bingo

OpenStudy (rootbeer003):

Nick is solving the equation 3x^2 = 20 – 7x with the quadratic formula. What values should he use for a, b, and c? a = 3 b = 20 c = –7 a = 3 b = –7 c = 20 a = 3 b = 7 c = –20 a = 3 b = –20 c = 7

OpenStudy (rootbeer003):

c?

OpenStudy (anonymous):

a trick question

OpenStudy (anonymous):

ok that was wrong !

TheSmartOne (thesmartone):

3x^2 = 20 – 7x 3x^2 + 7x - 20 = 0 a = 3 b = 7 c = -20

OpenStudy (anonymous):

\[3x^2+7x-20=0\]is the first step

OpenStudy (anonymous):

so yeah, C

TheSmartOne (thesmartone):

and that's in the form of \(ax^2 +bx + c\)

OpenStudy (rootbeer003):

really

TheSmartOne (thesmartone):

yes

OpenStudy (anonymous):

i know it was a totally lucky guess

OpenStudy (rootbeer003):

yes it was

OpenStudy (rootbeer003):

ok next

OpenStudy (rootbeer003):

Calculate the discriminant. x^2 + 2x – 2 = 0 –8 –4 8 12

OpenStudy (rootbeer003):

b?

TheSmartOne (thesmartone):

discriminant = b^2 - 4ac 2^2 - 4(1)(2) 4 - 8 -4

TheSmartOne (thesmartone):

so correct

OpenStudy (anonymous):

damn ! another lucky guess or did you do it?

OpenStudy (rootbeer003):

i did it

OpenStudy (rootbeer003):

What number would you have to add to both sides to complete the square? x^2 – 16x = –8 –16 –64 16 64

OpenStudy (rootbeer003):

c?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

half of 16 is ?

TheSmartOne (thesmartone):

half of 16?

OpenStudy (rootbeer003):

8

TheSmartOne (thesmartone):

and then square that ;p

OpenStudy (rootbeer003):

d

OpenStudy (anonymous):

and then square that ;p

OpenStudy (rootbeer003):

wow nice

OpenStudy (anonymous):

yeah D

TheSmartOne (thesmartone):

yeah D

OpenStudy (rootbeer003):

Solve by using the perfect squares method. x^2 + 12x + 36 = 0 –18 18 6 –6

OpenStudy (rootbeer003):

-6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[(x+6)^2=0\\ x+6=0\\ x=-6\]

TheSmartOne (thesmartone):

(x+6)^2 = 0 x = -6

OpenStudy (rootbeer003):

Lijuan is solving the equation x^2 – x = 14 with the quadratic formula. What values should she use for a, b, and c? a = –1, b = 1, c = 14 a = 1, b = –1, c = 14 a = –1, b = 1, c = –14 a = 1, b = –1, c = –14

OpenStudy (anonymous):

got another lucky guess?

OpenStudy (rootbeer003):

hahahahahhahahahaha no

OpenStudy (anonymous):

oh

TheSmartOne (thesmartone):

x^2 – x = 14 x^2 - x - 14 = 0 ax^2 - bx - c = 0 what is a, b, and c?

OpenStudy (anonymous):

\[x^2-x-14=0\] how about now?

OpenStudy (rootbeer003):

c?

OpenStudy (anonymous):

no, a is not minus one

OpenStudy (rootbeer003):

d

OpenStudy (anonymous):

yes

OpenStudy (rootbeer003):

Use the discriminant to determine the number and type of solutions the equation has. x^2 + 6x + 12 = 0 two irrational solutions no real solution one real solution two rational solutions

TheSmartOne (thesmartone):

b^2 - 4ac 6^2 - 4(1)(12) simplify it more

OpenStudy (anonymous):

wonder why they ask to do this by the discriminant always

OpenStudy (rootbeer003):

a?

OpenStudy (anonymous):

\[x^2 + 6x + 12 = 0 \\ x^2+6x=-12\\ (x+3)^2=-12+9=-3\] nope a square can't be negative

TheSmartOne (thesmartone):

bingo, it's a

OpenStudy (anonymous):

no it aint

OpenStudy (anonymous):

\[36-48=-12\] is negative no real solution

TheSmartOne (thesmartone):

my bad, got confused with imaginary and irrational

OpenStudy (rootbeer003):

this should be my pfp

OpenStudy (anonymous):

yeah you'd get lots of help then...